the flow in a lake is 0.5 m ^ 3 / s. a water fall in the lake has a height of 1.5m. with a thermal efficiency of 40%, how much power can be added


1
Expert's answer
2021-04-30T11:27:05-0400

"P=\eta P_0=\frac {\eta A}{t}=\frac{\eta Fh}{t}=\frac{\eta mgh}{t}=\frac{\eta \rho Vgh}{t}=\eta \rho Qgh=0.4\cdot1000\cdot 0.5\cdot 10\cdot 1.5=3~kW."